If your experiment, you compared the scores of a set of 20 individuals before and after taking a drug. Your obtained statistic is = 1.58 You re interested in finding a statistically significant difference. Alpha level is 0.05. What is the critical value?
Answer #1

This is two tailed test.

df = n - 1 = 20 - 1 = 19

T critical values at 0.05 significance level with 19 df = -2.093 , 2.093